1. If 18.8 m/ of 1.25 M sodium hydroxide solution is necessary to netralized 25.0 m/ of a hydrochloric acid solution to a phenolphtalein end point, calculate the molarity and the percent by mass (density of solution= 1.02 g/ m/) of the hydrochloric acid solution.
1
Expert's answer
2022-05-16T04:12:03-0400

NaOH + HCl = NaCl + H2O


n(NaOH) = C(NaOH)*V(NaOH) = 1.25 M * 0.0188 L = 0.0235 moles

n(HCl) = n(NaOH) = 0.0235 moles


C(HCl) = n(HCl) / V(HCl) = 0.0235 moles / 0.025 L = 0.94 M


ω(HCl) = (C(HCl) * Mr(HCl) / (ρ(HCl)*1000)) * 100% = (0.94 M * 36.5 g/mol) / (1.02 g/ml * 1000) = 0.0336 * 100% = 3.36 %